There are 480 dishes that need to be rinsed. John can rinse them in 80 minutes by himself. It will take his friend Bob 120 minutes to rinse these dishes. How long will it take them if they rinse these 480 dishes together?

Answer by checkley77(12569)

1/80+1/120=1/x
(120+80)/80*120=1/x
200/9,600=1/x cross multiply
200x=9,600
x=9,600/200
x=48 minutes working together.